# Federated Learning for Smart Grid: A Survey on Applications and Potential Vulnerabilities

### v1

@@ 4.2.3. Voltage Control
Voltage regulation is a critical process in electrical power systems, designed to maintain voltage levels within acceptable limits across the network. It involves controlling and adjusting the voltage to ensure stable and reliable power delivery to consumers (Xie et al., 2021). However, these networks are susceptible to data-related attacks, which can lead to voltage instability. Attackers may exploit vulnerabilities in communication channels to inject false data or disrupt the flow of information critical for effective voltage control.
In the FL-based SG system, Liu et al. (Liu and Wu, 2022) combine FL with reinforcement learning techniques to develop a privacy-preserving and efficient control strategy. In this federated reinforcement learning framework, grid-connected interface inverters of renewable energy sources are modeled as deep neural network-based agents that interact with their local environment to learn optimal voltage-var control strategies. The process operates in a two-stage process: local training using reinforcement learning algorithms, followed by periodic global aggregation of model updates. This approach offers several advantages, including enhanced data privacy, reduced communication costs, improved scalability, and adaptability to changing network conditions.
@@ 4.3. Vulnerabilities in Consumption Stage
The consumption part of a SG represents the end-user side of the energy distribution system (Kaur et al., 2020), encompassing homes, businesses, industries, and more. It integrates advanced technologies like AMI with smart meters for real-time usage monitoring and demand-side management to optimize energy consumption patterns. This system leverages IoT devices and real-time analytics to enhance energy efficiency and provide consumers with detailed insights into their usage. Additionally, sophisticated forecasting techniques and machine learning algorithms are employed to predict and manage energy loads effectively, ensuring a balance between supply and demand. However, smart meters and IoT devices, which are integral to these systems, can be vulnerable to attacks (Liu et al., 2021), potentially allowing malicious customers to manipulate energy consumption data or disrupt power supply.
@@ 4.3.1. Load Forecasting
Smart meter measurements, while essential for accurate demand forecasting, raise significant privacy and data breach concerns for consumers. The fine-grained load data collected by smart meters can potentially reveal sensitive information about residential users’ daily routines and behaviors, posing risks to property security (Husnoo et al., 2023). In the FL framework, Husnoo et al. (Husnoo et al., 2023) propose a secure FL approach designed for residential short-term load forecasting. This approach addresses the critical balance between accurate demand prediction and protecting consumer privacy in smart meter data. FL allows collaborative model training without exposing raw data, as models are trained locally, and only parameters are shared with a central server. To enhance security, the framework incorporates advanced techniques like gradient quantization using SignSGD (Bernstein et al., 2018) and differential privacy (Dwork, 2006). Additionally, it employs hierarchical and peer-to-peer strategies to improve both privacy and accuracy. The system tackles challenges such as non-IID data distribution and potential attacks through methods like deep federated adaptation and secure aggregation. Experimental results demonstrate that this approach achieves high forecasting accuracy while preserving user data privacy, with significant improvements in prediction accuracy compared to traditional methods.
@@ 4.3.2. Energy Theft Detection
Energy theft detection (Salinas and Li, 2015) in SGs refers to the process of identifying and preventing unauthorized consumption of electricity, an illegal activity that causes significant economic losses to utility companies. SGs integrate digital technologies with traditional power systems, enabling advanced energy management and monitoring capabilities. These systems utilize AMI and smart meters to collect, measure, and analyze energy usage data from customers at regular intervals. However, the collection and analysis of consumer energy data create vulnerabilities, potentially allowing attackers to compromise smart meters.
Within the FL framework, Wen et al. (Wen et al., 2021) introduce FedDetect, a novel privacy-preserving FL framework designed for energy theft detection in SGs. FedDetect leverages FL to train machine learning models on distributed devices, aggregating local model updates instead of raw data to protect user privacy. Ashraf et al. (Ashraf et al., 2022) propose FedDP, a privacy-protecting theft detection scheme for SGs using FL. This framework addresses the challenge of utilizing consumer energy data for theft detection while maintaining privacy protection. Zafar et al. (Zafar et al., 2023) develop a FL-assisted hybrid deep learning model for electricity theft detection in SGs. Their model combines FL techniques with deep learning to enhance security and reliability of electricity theft detection. Li et al. (Li et al., 2024a) present FUSE, a novel electricity theft detection framework that integrates FL with U-shape split learning. FUSE aims to address the challenges of data privacy and security in SG systems while effectively detecting electricity theft.
@@ 5.1.1. Byzantine Attack in FL
Min-Max/Min-Sum attack (Shejwalkar and Houmansadr, 2021). The creation of malicious model updates involves two steps. First, the attacker generates a malicious update by perturbing the average of all benign updates. In the second step, specifically for the Min-Max attack, the attacker optimizes this malicious update to ensure that its maximum Euclidean distance from any benign update is no greater than the maximum distance between any two benign updates. As for Min-Sum attack, the malicious update is optimized to ensure that the sum of its distances with each benign update is upper-bounded by the maximum total distance of a benign update among other benign updates. The Min-Max and Min-Sum attacks are powerful since malicious clients try to search for an optimal perturbation such that it has maximum impact.
AGR-tailored Trimmed-mean attack (Shejwalkar and Houmansadr, 2021). In (Shejwalkar and Houmansadr, 2021), the authors proposed several attack methods specific to some well-known defense methods. We take the AGR-tailored Trimmed-mean attack as a example. The AGR-tailored Trimmed-mean is designed to attack the defense method Trimmed-mean proposed in (Yin et al., 2018a) by maximizing the Euclidean distance between the aggregated result of simple average and Trmean. This straightforward and intuitive method is effective, but the malicious model updates produced by the AGR-tailored Trimmed-mean attack differ significantly from benign ones, making them susceptible to easy detection by defense methods.
Lie attack (Baruch et al., 2019). In A-little-is-enough attack, the malicious clients apply slight perturbation to their local benign updates, making it hard to be detected by the defense methods. Specifically, the malicious clients calculate the element-wise mean $\mu_{j}$ and standard error $\sigma_{j}$ of all updates and generate the element of malicious updates by $(\beta_{i})_{j}=\mu_{j}-z\times\sigma_{j}$, where $\beta$ is the malicious model updates and $j\in[d]$. Here, $z$ is a scaling factor decided by the number of malicious clients and the total number of clients.
ByzMean attack (Xu et al., 2022). The ByzMean attack manipulates the mean of updates to be an arbitrarily malicious update. Specifically, it divides the malicious clients into two groups, with $m_{1}$ and $m_{2}$ clients respectively. The first group uses any existing attack methods (e.g., Lie attack) to generate their malicious updates. The second group creates their malicious updates to ensure that the overall average of all updates matches the average of the malicious updates from the first group.
@@ 5.1.2. Backdoor Attack in FL
Backdoor attacks (aka data poisoning attacks) in FL aim to maintain the global model’s performance on clean inputs while causing it to make incorrect predictions on inputs containing a specific predefined feature (also called trigger). Since backdoor attacks (Lyu et al., 2024, 2023) do not significantly affect the main task accuracy, the malicious local model updates are similar to benign ones, making anomaly detection more challenging in FL. We introduce several classical and commonly discussed backdoor attacks as follows.
Badnet attack (Gu et al., 2017). Badnet is a simple and effective backdoor attack where the attacker injects the same trigger into the training data of compromised clients. The ground truth label of these training data will also be changed to another backdoor class.
Distributed backdoor attack (Xie et al., 2019). In the distributed backdoor attack (DBA), the adversary decomposes the backdoor trigger into several small triggers and applies them to different compromised clients rather than using the same trigger for all malicious clients (Badnet attack). This makes the attack harder to detect, as each client’s updates appear less suspicious.
Scaling attack (Bagdasaryan et al., 2020). Based on Badnet attack, a scaling attack involves malicious clients scaling their model updates in a way that scales up the impact of their backdoor-triggered updates without affecting the clean data performance. The scaling factor ensures that the backdoor pattern has a stronger influence on the global model.
PGD attack (Wang et al., 2020a). The Projected Gradient Descent (PGD) attack is adapted from adversarial machine learning. It iteratively perturbs the model updates of compromised clients within a certain bound to embed the backdoor. This method is effective at creating small but impactful changes to the model parameters.
Neurotoxin attack (Zhang et al., 2022). Neurotoxin explores a persistent attack method designed for scenarios where attackers can only participate in a limited number of rounds. Specifically, in the Neurotoxin attack, the malicious clients project their model updates into a subspace that excludes the global coordinates. This approach ensures that the projected updates from malicious clients are primarily embedded in coordinates with minimal perturbation from benign updates, maintaining the backdoor effect even after the attack ceases.
@@ 5.1.3. Inference Attack in FL
Membership Inference Attack (MIA) (Hayes et al., 2019; Shokri et al., 2017) is a major privacy attack that seeks to determine whether a specific data record was part of the training dataset used to build a machine learning model. In the context of FL, MIA poses a significant threat to the privacy of participating clients and their data. Various entities, including malicious clients, servers, or external attackers, can execute MIA in FL. These adversaries exploit the model’s natural tendency to overfit on training data, leading to unintended information leakage. Approaches to conducting MIA in FL include attack models (Nasr et al., 2019) and gradient-based MIAs, which exploit the gradients exchanged during the FL process. Passive MIA techniques also exist, where adversaries infer membership information by merely observing communication between clients and the server. Successful MIAs compromise the privacy of individual users, a particularly serious issue in sensitive domains such as healthcare or finance, where data confidentiality is paramount. Beyond identifying membership, MIA can also reveal attributes or properties of the training data, further threatening user privacy. Defending against MIA in FL presents unique challenges due to the distributed nature of the learning process, with traditional privacy-preserving techniques either being inapplicable or significantly impacting model utility. As a result, researchers are actively developing novel defense mechanisms to improve FL’s resilience to MIA while preserving model performance.
@@ 5.2. SOTA FL Defense
To defend against the attacks mentioned above, the server in FL usually adopts a defense method. Existing defense methods can be generally divided into three categories: detection-filtering methods (Blanchard et al., 2017; Guerraoui et al., 2018; Xu and Huang, 2022; Xu et al., 2024, 2022; Shejwalkar and Houmansadr, 2021; Cao et al., 2020), impact-reduction methods(Yin et al., 2018a; Guerraoui et al., 2018; Fung et al., 2020; Huang et al., 2024; Pillutla et al., 2022; Panda et al., 2022; Zhang and Hu, 2023), and differential privacy-based methods.
@@ 5.2.1. Detection-Filtering Methods
Detection-filtering methods aim to maximize robustness against attacks by identifying and removing malicious local model updates from the aggregation process. These methods typically rely on one or more statistical metrics (e.g., norms, cosine similarity, Euclidean distance) to assess local model updates. By comparing the extracted metrics with each update, they filter out updates that exhibit abnormal metric values, preventing them from influencing the global model. For example, Krum (Blanchard et al., 2017) selects the most reliable local model that has the smallest sum of squared Euclidean distances to all other models as the output. Multi-Krum (Blanchard et al., 2017) extends this by selecting multiple reliable models for averaging instead of just one. In (Xu and Huang, 2022), the authors applied the KL-divergence to measure the distance between the empirical sign distributions of the local models in two consecutive rounds and used it to employ a weighted average of models for each client. The Divide-and-Conquer (DnC) method proposed in (Shejwalkar and Houmansadr, 2021) calculates the product of a model with its principle component and removes the models with large projection values. FLTrust (Cao et al., 2020) leverages a root dataset and trains a root global model based on it. After collecting local updates, FLTrust aggregates those updates with high Cosine similarity to root model updates. SignGuard (Xu et al., 2022) combines direction-based clustering and magnitude-based filtering to identify malicious model updates, where the direction and magnitude are obtained by the sign statistics and $L_{2}$-norm of the model update, respectively. Recently, a defense method called LASA that achieves SOTA results has been proposed in (Xu et al., 2024). It employs a layer-wise adaptive filter to adaptively select benign sparsified layers using both magnitude and direction metrics across all clients for aggregation.
@@ 5.2.2. Impact-Reduction Methods
Impact-reduction methods try to integrate all model updates but employ strategies to reduce the impact of malicious updates. For instance, Trimmed mean (Trmean) proposed in (Yin et al., 2018a) discards a certain percentage of the highest and lowest values among the received models from clients for each dimension. After this trimming, the mean of the remaining values is computed by the server, which mitigates the impact of extreme values on the aggregated model. Mhamdi et al. (Guerraoui et al., 2018) proposed Bulyan that combines Trmean and Multi-Krum to extract benign models. In (Pillutla et al., 2022), the authors proposed to use the geometric median (GeoMed) of all local models as the aggregated model, which is effective in filtering out the outlier values. RLR (Ozdayi et al., 2021) uses the sign information of model updates to adjust the global learning rate for each model coordinate without distinguishing between benign and malicious updates, resulting in poor accuracy. Foolsgold (Fung et al., 2020) assumes that the malicious updates are consistent with each other. It assigns aggregation weights to model updates based on the maximum Cosine similarity between the last layers of pairwise model updates. A higher Cosine similarity value indicates a higher probability that the updates are malicious, leading to smaller aggregation weights being assigned. SparseFed proposed in (Panda et al., 2022) sparsifies the aggregated model update at the server side, integrating with model clipping and error feedback, to mitigate the impact of malicious local model updates. In addition, FedVRDP proposed in (Zhang and Hu, 2023) applies sparsification on local model updates using a mask obtained from the global model in the previous round and combines it with random Gaussian noise perturbation and variance reduction techniques, to achieve robustness and differential privacy protection simultaneously. A recent approach called Lockdown (Huang et al., 2024) assumes that those malicious parameters served to recognize backdoor triggers will be deemed unimportant for benign clients and applies sparsification to prune unimportant parameters on the local model.

### v2

@@ 4.2.3. Voltage Control
Voltage regulation is a critical process in electrical power systems, designed to maintain voltage levels within acceptable limits across the network. It involves controlling and adjusting the voltage to ensure stable and reliable power delivery to consumers (xie2021review). However, these networks are susceptible to data-related attacks, which can lead to voltage instability. Attackers may exploit vulnerabilities in communication channels to inject false data or disrupt the flow of information critical for effective voltage control.
In the FL-based SG system, Liu et al. (liu2022federated) combine FL with reinforcement learning techniques to develop a privacy-preserving and efficient control strategy. In this federated reinforcement learning framework, grid-connected interface inverters of renewable energy sources are modeled as deep neural network-based agents that interact with their local environment to learn optimal voltage-var control strategies. The process operates in a two-stage process: local training using reinforcement learning algorithms, followed by periodic global aggregation of model updates. This approach offers several advantages, including enhanced data privacy, reduced communication costs, improved scalability, and adaptability to changing network conditions.
@@ 4.3. Vulnerabilities in Consumption Stage
The consumption part of a SG represents the end-user side of the energy distribution system (kaur2020energy), encompassing homes, businesses, industries, and more. It integrates advanced technologies like AMI with smart meters for real-time usage monitoring and demand-side management to optimize energy consumption patterns. This system leverages IoT devices and real-time analytics to enhance energy efficiency and provide consumers with detailed insights into their usage. Additionally, sophisticated forecasting techniques and machine learning algorithms are employed to predict and manage energy loads effectively, ensuring a balance between supply and demand. However, smart meters and IoT devices, which are integral to these systems, can be vulnerable to attacks (liu2021vulnerability), potentially allowing malicious customers to manipulate energy consumption data or disrupt power supply.
@@ 4.3.1. Load Forecasting
Smart meter measurements, while essential for accurate demand forecasting, raise significant privacy and data breach concerns for consumers. The fine-grained load data collected by smart meters can potentially reveal sensitive information about residential users’ daily routines and behaviors, posing risks to property security (husnoo2023secure). In the FL framework, Husnoo et al. (husnoo2023secure) propose a secure FL approach designed for residential short-term load forecasting. This approach addresses the critical balance between accurate demand prediction and protecting consumer privacy in smart meter data. FL allows collaborative model training without exposing raw data, as models are trained locally, and only parameters are shared with a central server. To enhance security, the framework incorporates advanced techniques like gradient quantization using SignSGD (bernstein2018signsgd) and differential privacy (dwork2006differential). Additionally, it employs hierarchical and peer-to-peer strategies to improve both privacy and accuracy. The system tackles challenges such as non-IID data distribution and potential attacks through methods like deep federated adaptation and secure aggregation. Experimental results demonstrate that this approach achieves high forecasting accuracy while preserving user data privacy, with significant improvements in prediction accuracy compared to traditional methods.
@@ 4.3.2. Energy Theft Detection
Energy theft detection (salinas2015privacy) in SGs refers to the process of identifying and preventing unauthorized consumption of electricity, an illegal activity that causes significant economic losses to utility companies. SGs integrate digital technologies with traditional power systems, enabling advanced energy management and monitoring capabilities. These systems utilize AMI and smart meters to collect, measure, and analyze energy usage data from customers at regular intervals. However, the collection and analysis of consumer energy data create vulnerabilities, potentially allowing attackers to compromise smart meters.
Within the FL framework, Wen et al. (wen2021feddetect) introduce FedDetect, a novel privacy-preserving FL framework designed for energy theft detection in SGs. FedDetect leverages FL to train machine learning models on distributed devices, aggregating local model updates instead of raw data to protect user privacy. Ashraf et al. (ashraf2022feddp) propose FedDP, a privacy-protecting theft detection scheme for SGs using FL. This framework addresses the challenge of utilizing consumer energy data for theft detection while maintaining privacy protection. Zafar et al. (zafar2023step) develop a FL-assisted hybrid deep learning model for electricity theft detection in SGs. Their model combines FL techniques with deep learning to enhance security and reliability of electricity theft detection. Li et al. (li2024fuse) present FUSE, a novel electricity theft detection framework that integrates FL with U-shape split learning. FUSE aims to address the challenges of data privacy and security in SG systems while effectively detecting electricity theft.
@@ 5.1.1. Byzantine Attack in FL
Min-Max/Min-Sum attack (dnc). The creation of malicious model updates involves two steps. First, the attacker generates a malicious update by perturbing the average of all benign updates. In the second step, specifically for the Min-Max attack, the attacker optimizes this malicious update to ensure that its maximum Euclidean distance from any benign update is no greater than the maximum distance between any two benign updates. As for Min-Sum attack, the malicious update is optimized to ensure that the sum of its distances with each benign update is upper-bounded by the maximum total distance of a benign update among other benign updates. The Min-Max and Min-Sum attacks are powerful since malicious clients try to search for an optimal perturbation such that it has maximum impact.
AGR-tailored Trimmed-mean attack (dnc). In (dnc), the authors proposed several attack methods specific to some well-known defense methods. We take the AGR-tailored Trimmed-mean attack as a example. The AGR-tailored Trimmed-mean is designed to attack the defense method Trimmed-mean proposed in (trmean) by maximizing the Euclidean distance between the aggregated result of simple average and Trmean. This straightforward and intuitive method is effective, but the malicious model updates produced by the AGR-tailored Trimmed-mean attack differ significantly from benign ones, making them susceptible to easy detection by defense methods.
Lie attack (lie). In A-little-is-enough attack, the malicious clients apply slight perturbation to their local benign updates, making it hard to be detected by the defense methods. Specifically, the malicious clients calculate the element-wise mean $\mu_{j}$ and standard error $\sigma_{j}$ of all updates and generate the element of malicious updates by $(\beta_{i})_{j}=\mu_{j}-z\times\sigma_{j}$, where $\beta$ is the malicious model updates and $j\in[d]$. Here, $z$ is a scaling factor decided by the number of malicious clients and the total number of clients.
ByzMean attack (signguard). The ByzMean attack manipulates the mean of updates to be an arbitrarily malicious update. Specifically, it divides the malicious clients into two groups, with $m_{1}$ and $m_{2}$ clients respectively. The first group uses any existing attack methods (e.g., Lie attack) to generate their malicious updates. The second group creates their malicious updates to ensure that the overall average of all updates matches the average of the malicious updates from the first group.
@@ 5.1.2. Backdoor Attack in FL
Backdoor attacks (aka data poisoning attacks) in FL aim to maintain the global model’s performance on clean inputs while causing it to make incorrect predictions on inputs containing a specific predefined feature (also called trigger). Since backdoor attacks (lyu2024task; lyu2023attention) do not significantly affect the main task accuracy, the malicious local model updates are similar to benign ones, making anomaly detection more challenging in FL. We introduce several classical and commonly discussed backdoor attacks as follows.
Badnet attack (badnet). Badnet is a simple and effective backdoor attack where the attacker injects the same trigger into the training data of compromised clients. The ground truth label of these training data will also be changed to another backdoor class.
Distributed backdoor attack (dba). In the distributed backdoor attack (DBA), the adversary decomposes the backdoor trigger into several small triggers and applies them to different compromised clients rather than using the same trigger for all malicious clients (Badnet attack). This makes the attack harder to detect, as each client’s updates appear less suspicious.
Scaling attack (scaling). Based on Badnet attack, a scaling attack involves malicious clients scaling their model updates in a way that scales up the impact of their backdoor-triggered updates without affecting the clean data performance. The scaling factor ensures that the backdoor pattern has a stronger influence on the global model.
PGD attack (PGD). The Projected Gradient Descent (PGD) attack is adapted from adversarial machine learning. It iteratively perturbs the model updates of compromised clients within a certain bound to embed the backdoor. This method is effective at creating small but impactful changes to the model parameters.
Neurotoxin attack (Neurotoxin). Neurotoxin explores a persistent attack method designed for scenarios where attackers can only participate in a limited number of rounds. Specifically, in the Neurotoxin attack, the malicious clients project their model updates into a subspace that excludes the global coordinates. This approach ensures that the projected updates from malicious clients are primarily embedded in coordinates with minimal perturbation from benign updates, maintaining the backdoor effect even after the attack ceases.
@@ 5.1.3. Inference Attack in FL
Membership Inference Attack (MIA) (membership1; membership2) is a major privacy attack that seeks to determine whether a specific data record was part of the training dataset used to build a machine learning model. In the context of FL, MIA poses a significant threat to the privacy of participating clients and their data. Various entities, including malicious clients, servers, or external attackers, can execute MIA in FL. These adversaries exploit the model’s natural tendency to overfit on training data, leading to unintended information leakage. Approaches to conducting MIA in FL include attack models (membership3) and gradient-based MIAs, which exploit the gradients exchanged during the FL process. Passive MIA techniques also exist, where adversaries infer membership information by merely observing communication between clients and the server. Successful MIAs compromise the privacy of individual users, a particularly serious issue in sensitive domains such as healthcare or finance, where data confidentiality is paramount. Beyond identifying membership, MIA can also reveal attributes or properties of the training data, further threatening user privacy. Defending against MIA in FL presents unique challenges due to the distributed nature of the learning process, with traditional privacy-preserving techniques either being inapplicable or significantly impacting model utility. As a result, researchers are actively developing novel defense mechanisms to improve FL’s resilience to MIA while preserving model performance.
@@ 5.2. SOTA FL Defense
To defend against the attacks mentioned above, the server in FL usually adopts a defense method. Existing defense methods can be generally divided into three categories: detection-filtering methods (krum; bulyan; xu2022byzantine; LASA; signguard; dnc; fltrust), impact-reduction methods(trmean; bulyan; Foolsgold; lockdown; geomed; sparsefed; zhang2023byzantine), and differential privacy-based methods.
@@ 5.2.1. Detection-Filtering Methods
Detection-filtering methods aim to maximize robustness against attacks by identifying and removing malicious local model updates from the aggregation process. These methods typically rely on one or more statistical metrics (e.g., norms, cosine similarity, Euclidean distance) to assess local model updates. By comparing the extracted metrics with each update, they filter out updates that exhibit abnormal metric values, preventing them from influencing the global model. For example, Krum (krum) selects the most reliable local model that has the smallest sum of squared Euclidean distances to all other models as the output. Multi-Krum (krum) extends this by selecting multiple reliable models for averaging instead of just one. In (xu2022byzantine), the authors applied the KL-divergence to measure the distance between the empirical sign distributions of the local models in two consecutive rounds and used it to employ a weighted average of models for each client. The Divide-and-Conquer (DnC) method proposed in (dnc) calculates the product of a model with its principle component and removes the models with large projection values. FLTrust (fltrust) leverages a root dataset and trains a root global model based on it. After collecting local updates, FLTrust aggregates those updates with high Cosine similarity to root model updates. SignGuard (signguard) combines direction-based clustering and magnitude-based filtering to identify malicious model updates, where the direction and magnitude are obtained by the sign statistics and $L_{2}$-norm of the model update, respectively. Recently, a defense method called LASA that achieves SOTA results has been proposed in (LASA). It employs a layer-wise adaptive filter to adaptively select benign sparsified layers using both magnitude and direction metrics across all clients for aggregation.
@@ 5.2.2. Impact-Reduction Methods
Impact-reduction methods try to integrate all model updates but employ strategies to reduce the impact of malicious updates. For instance, Trimmed mean (Trmean) proposed in (trmean) discards a certain percentage of the highest and lowest values among the received models from clients for each dimension. After this trimming, the mean of the remaining values is computed by the server, which mitigates the impact of extreme values on the aggregated model. Mhamdi et al. (bulyan) proposed Bulyan that combines Trmean and Multi-Krum to extract benign models. In (geomed), the authors proposed to use the geometric median (GeoMed) of all local models as the aggregated model, which is effective in filtering out the outlier values. RLR (rlr) uses the sign information of model updates to adjust the global learning rate for each model coordinate without distinguishing between benign and malicious updates, resulting in poor accuracy. Foolsgold (Foolsgold) assumes that the malicious updates are consistent with each other. It assigns aggregation weights to model updates based on the maximum Cosine similarity between the last layers of pairwise model updates. A higher Cosine similarity value indicates a higher probability that the updates are malicious, leading to smaller aggregation weights being assigned. SparseFed proposed in (sparsefed) sparsifies the aggregated model update at the server side, integrating with model clipping and error feedback, to mitigate the impact of malicious local model updates. In addition, FedVRDP proposed in (zhang2023byzantine) applies sparsification on local model updates using a mask obtained from the global model in the previous round and combines it with random Gaussian noise perturbation and variance reduction techniques, to achieve robustness and differential privacy protection simultaneously. A recent approach called Lockdown (lockdown) assumes that those malicious parameters served to recognize backdoor triggers will be deemed unimportant for benign clients and applies sparsification to prune unimportant parameters on the local model.